40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 18 12:27:42 2019, gautam, Update, General, IFO status IMG_7343.JPGITMXOL.png
    Reply  Mon Mar 18 17:06:12 2019, gautam, Update, Optical Levers, ITMY HeNe replaced OLRIN.pdfOL_PIT.pdfOL_YAW.pdf
    Reply  Tue May 28 17:41:13 2019, gautam, Update, General, IFO status IMG_7444.JPG
Message ID: 14642     Entry time: Tue May 28 17:41:13 2019     In reply to: 14483
Author: gautam 
Type: Update 
Category: General 
Subject: IFO status 

[chub, gautam]

Today, we tried to resuscitate the c1iscaux2 channels by swapping the existing, failed VME crate with the newly freed up crate from c1susaux. In summary, the crate gets power, and the EPICS server gets satrted, but I am unable to switch the whitening gain on the whitening boards. I belive that this has to do with the FAIL LEDs that are on for the XVME-220 units. We were careful to preserve the location of the various cards in the VME crates during the swap. Rather than do a detailed debugging with custom RJ45 cables and terminal emulators, I think we should just focus the efforts on getting the Acromag system up and running.

Our work must have bumped a cable to the c1lsc expansion chassis in the same rack - the c1lsc FE had crashed. I rebooted it using the script - everything came back gracefully.

Attachment 1: IMG_7444.JPG  1.728 MB  Uploaded Tue May 28 18:46:57 2019  | Hide | Hide all
IMG_7444.JPG
ELOG V3.1.3-